After getting a 15% discount on the
price of a new Chrysler Sebring convertible, Helen
paid $27,000. What was the original price of the
convertible to the nearest dollar?
: After getting a 15% discount on the
price of a new Chrysler Sebring convertible, Helen
paid $27,000. What was the original price of the
convertible to the nearest dollar?

Answer by oscargut(666)

let x the original price
then x-0.15x=27000
then 0.85x=27000
then x=27000/0.85=31764.70588 (aprox)
answer: original price is $31765